5x+15= 20 2x+6y = 12

solve for x and y
Answer #1

500 + 15 = 20 22 t by - 0) from ean i 5x = 20-15 5x = 5 1 x put in ega lii) 2(x) + 6ly) = 12 2x1 + 6y = 12 6y= 12:02 by = 10